Calculate the standard entropy change for the following using the table of standard values.(first, predict the sign for ∆S qualitatively)

2H2(g)+O2(g)→2H2O(g)


1
Expert's answer
2021-06-14T13:34:38-0400

Answer -:

At 298K as a standard temperature:

  • "S\degree"(H2) = 130.68 J/mol K
  • "S\degree"(O2) = 205.14 J/mol K
  • "S\degree"(H20) = 188.83 J/mol K


From the balanced equation we can write the equation for "S\degree"(the change in the standard molar entropy for the reaction):

"S\degree = 2*S\degree(H_2O) - [2*S\degree(H_2)+S\degree(O_2)]\\
S\degree=2*188.83 - [ 2*130.68+205.14]\\
S\degree= -88.84 J/mol K"

It would appear that the process result in a decrease in entropy -i.e. a decrease in disorder. This is expected because we are decreasing the number of gas molecules.